### v2

@@ Abstract
Ultrasound imaging of the medial elbow is crucial for the early diagnosis of Ulnar Collateral Ligament (UCL) injuries. Specifically, measuring the elbow joint space in ultrasound images is used to assess the valgus instability of the elbow caused by UCL injuries. To automate this measurement, a model trained on a precisely annotated dataset is necessary; however, no publicly available dataset exists to date. This study introduces a novel ultrasound medial elbow dataset to measure the joint space. The dataset comprises 4,201 medial elbow ultrasound images from 22 subjects, with landmark annotations on the humerus and ulna, based on the expertise of three orthopedic surgeons. We evaluated joint space measurement methods on our proposed dataset using heatmap-based, regression-based, and token-based landmark detection methods. While heatmap-based landmark detection methods generally achieve high accuracy, they sometimes produce multiple peaks on a heatmap, leading to incorrect detection. To mitigate this issue and enhance landmark localization, we propose Shape Subspace (SS) landmark refinement by measuring geometrical similarities between the detected and reference landmark positions. The results show that the mean joint space measurement error is 0.116 mm when using HRNet. Furthermore, SS landmark refinement can reduce the mean absolute error of landmark positions by 0.010 mm with HRNet and by 0.103 mm with ViTPose on average. These highlight the potential for high-precision, real-time diagnosis of UCL injuries by accurately measuring joint space. Lastly, we demonstrate point-based segmentation for the humerus and ulna using the detected landmarks as inputs. Our dataset is publicly available at https://github.com/Akahori000/Ultrasound-Medial-Elbow-Dataset.
keywords:
@@
Ultrasound Dataset, Medial Elbow, Landmark Detection, UCL injury Diagnosis, Shape Subspace
1 Background
1.1 Introduction
Ultrasound imaging is an accessible, cost-effective, and non-invasive diagnostic technique in clinical practice. In orthopedics, medial elbow ultrasound screening is commonly performed on athletes. Several studies have demonstrated that measuring the ulnohumeral joint space via ultrasound offers critical insights into the risk of Ulnar Collateral Ligament (UCL) injury [1, 2]. Specifically, an increased joint space under stress conditions indicates joint laxity and potential UCL damage. Shanley et al. [2] measured joint space by defining landmarks on the edges of the ulna and humerus in ultrasound medial elbow images. They demonstrated that pitchers with UCL injuries exhibit greater joint space compared to those without injuries ($6.5\pm 1.2$ mm vs $5.3\pm 1.2$ mm). Michinobu et al.[3] discussed optimal elbow positioning during ultrasonography to effectively diagnose joint laxity.
Several studies [4, 5, 6] have explored deep learning-based diagnosis using elbow ultrasound images. These have focused on osteochondritis dissecans or median nerve detection. However, none of the studies have discussed the automation of joint space measurement for diagnosing UCL injuries. While precise and consistent measurement is required, it is challenging due to anatomical variability, noise, and artifacts. Fig. 1(a) shows ultrasound images of the medial elbow from different participants, and Fig. 1(b) provides enlarged views of the humerus and ulna near the joint space. These images show that bone shapes vary in individuals, and bone edges are often indistinct. Also, hyperechoic lines and artifacts often appear near bone surfaces or within the joint space.
@@ 1.1 Introduction
Figure 1: (a) The ultrasound images of the medial elbow obtained from different participants (b) The enlarged views of the joint space positions. The left and right hyperechoic lines located in the middle of each image are the humerus and ulna, respectively. The space between humerus and ulna is the joint space to be measured.
@@
In this paper, we introduce a medial elbow ultrasound dataset consisting of 4,201 images from 22 individuals. The landmark annotation is made on the humerus and ulna based on the expertise of three orthopedic surgeons. We evaluated joint space measurement methods on our proposed dataset using several landmark detection approaches. While heatmap-based landmark detection methods generally achieve high accuracy, they sometimes produce multiple peaks on a heatmap, leading to incorrect detection. To mitigate this, we propose Shape Subspace landmark refinement by measuring the geometric differences between detected and reference landmark positions. The results show that our measurement method is sufficiently accurate to diagnose UCL injuries. Finally, we demonstrate point-based segmentation of the humerus and ulna using detected landmarks as inputs. , showing an application of our landmark detection in the medial elbow.
1.2 Related works
1.2.1 Landmark detection
[…]
@@ 1.2.1 Landmark detection
landmark Detection Methods Landmark detection, known as keypoint detection, is a method of locating keypoints in images. Various methods have been proposed for pose estimation.
Regression-based approaches [7, 8, 9, 10, 11, 12, 13] directly predict landmark coordinates. While these methods are effective in simplicity and speed, they often struggle with precision in complex poses.
@@
Heatmap-based methods overcome this by generating heatmaps where the intensity represents the likelihood of a landmark’s location, achieving better spatial accuracy. Various studies have employed Convolutional Neural Network (CNN) including [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Among these, U-Net [24] and its extensions are widely used in localization tasks of medical images. The architecture utilizes skip connections to copy the feature maps from each stage of the encoder and directly pass them to the decoder’s corresponding stage. HRNet [25] maintains high-resolution representations throughout the whole process, allowing for accurate localization with multi-resolution features. Recently, Transformer-based approaches [26, 27, 28, 29] have gained attention in pose estimation tasks, and these models leverage the self-attention mechanism to capture long-range dependencies to enhance precision and robustness. ViTPose [29] has a simple architecture consisting of plain Vision Transformers. Lastly, the token-based method [30, 31] learns the relationships between skeletal joints using tokens. PCT [31] designs each token to represent a substructure of joints.
[…]
@@
Landmark Detection in Ultrasound Images Several landmark detection approaches have been applied to ultrasound images to measure anatomical structures [32, 33, 34, 35, 36]. U-Net and its extensions are used to measure left ventricle dimensions [33, 35]. Mask R-CNN-based architecture has been employed to evaluate developmental dysplasia of the hip from infant ultrasound images [34]. HRNet-based regression architecture is proposed to assess infant growth [36].
1.2.2 Shape Subspace
@@
Shape Subspace (SS) is a method to represent the 3D points of a 3D structure as a linear subspace within a high-dimensional vector space. By measuring the similarity between two shape subspaces, the geometrical difference of 3D structures can be measured [37]. Shape subspace has been utilized to classify 3D objects and human faces [38, 39].
1.2.3 Point-based Segmentation
@@
2 Method
2.1 Novel Ultrasound Elbow Dataset
2.1.1 Data Collection
@@
We constructed an ultrasound medial elbow dataset of 4,201 images from 22 participants (ages 22–56, men and women, from five countries). The images were acquired using musculoskeletal ultrasonography (SONIMAGE MX1, KONICA MINOLTA, Tokyo, Japan) with an 11-MHz linear probe by two experienced orthopedic surgeons. The participants lay on a bed with their elbows flexed at 90 degrees following [3]. Then, a surgeon moved the gel-applied probe around the medial elbow, as shown in Fig. 2(c). The images capture the medial epicondyle, UCL, medial surface of the humeral trochlea, and coronoid process of the ulna, a key region for diagnosing medial elbow conditions [3]. The image resolution is 528 × 672 pixels, corresponding to 0.0567 mm per pixel. In the images, the humerus is positioned on the left and the ulna on the right. Fig. 1 shows sample images of different participants with the enlarged joint space views.
Figure 2: (a) The annotated landmarks: 1) Distal end of humeral trochlea, 2) Proximal end of ulnar coronoid process. Landmark 1 and 2 are the joint space measurement points. 3) Tip of humeral medial epicondyle, 4) Deep end of the intersection of humeral medial epicondyle and ulnar collateral ligament, 5) Distal end of humeral medial epicondyle, 6) Proximal point adjacent to the landmark 1 on the center of hyperechoic line of humeral trochlea, 7) Distal point adjacent to the landmark 2 on the center of hyperechoic line of ulnar coronoid process, 8) Distal point of ulnar coronoid process. The white line where landmarks 1, and 3 through 6 are located is the humerus, while the line where landmarks 2, 7, and 8 are located is the ulna. (b) The enlarged view of joint space. The yellow arrow is the joint space length to be measured. Landmark 1 and 2 are the joint space measurement points. (c) The participant’s posture during the examination.
2.1.2 Annotation
@@
The obtained 4,201 images were manually annotated with landmarks at sub-pixel levels by the authors under the supervision of the three orthopedic surgeons. The landmarks include two joint space measurement points: the distal end of the humeral trochlea, and the proximal end of the ulnar coronoid process, corresponding to the positions described in [2]. In addition to the joint space measurement points, six additional points on the humerus and ulna were annotated. These points are of particular interest to sonographers, who primarily focus on bone structures during ultrasound image interpretation. The annotated landmarks were meticulously reviewed and discussed by the three orthopedic experts, ensuring the reliability of the dataset.
@@
Fig. 2(a) shows the annotated landmarks: 1) Distal end of humeral trochlea, 2) Proximal end of ulnar coronoid process. Landmark 1 and 2 are the joint space measurement points. 3) Tip of humeral medial epicondyle, 4) Deep end of the intersection of humeral medial epicondyle and ulnar collateral ligament, 5) Distal end of humeral medial epicondyle, 6) Proximal point adjacent to the landmark 1 on the center of hyperechoic line of humeral trochlea, 7) Distal point adjacent to the landmark 2 on the center of hyperechoic line of ulnar coronoid process, 8) Distal point of ulnar coronoid process. The yellow line segment between landmark 1 and 2 in Fig. 2(b) shows the joint space to be measured.
2.2 Landmark Detection Models
@@
We compare several landmark detection models: HRNet [25], ViTPose [26], and U-Net [24] as heatmap-based methods, YOLOv8 [13] as a regression-based method, and PCT [31] as a token-based method.
[…]
@@
HRNet, ViTPose, and U-Net The landmark detection process of HRNet, ViTPose, and U-Net is represented in Equation 1, where $f$ is the detection model and $\mathbf{x}\in\mathbb{R}^{W\times H\times 3}$ is the input image with width $W$, height $H$, and 3 channels. $f(\mathbf{x})$ outputs heatmaps of size $W^{\prime}\times H^{\prime}\times K$, where $W^{\prime}$ and $H^{\prime}$ are the heatmap dimensions, and $K$ is the number of landmarks. The $\arg\max(f(\mathbf{x}))$ operation returns the rough landmark coordinates $\mathbf{p}^{\prime}\in\mathbb{R}^{K\times 2}$. The Unbiased Data Processing (UDP) function [41, 42] refines these to sub-pixel accuracy, yielding the final landmark positions $\mathbf{p}\in\mathbb{R}^{K\times 2}$, $(x,y)$ coordinates of K landmarks.
@@
[…]
PCT and YOLOv8 Equation 2 represents the landmark detection using YOLOv8 and PCT. $g$ is the landmark detection model. The output $g(\mathbf{x})\in\mathbb{R}^{K\times 2}$ directly predicts the coordinates of the $K$ landmarks.
@@
2.3 Shape Subspace Landmark Refinement
We propose using shape subspace for landmark refinement in heatmap-based landmark detection. While heatmap-based models are trained to produce a single peak on each heatmap, they may occasionally generate multiple or unclear peaks. In such cases, measuring the geometrical similarities between detected and reference landmark positions using shape subspace can help select the most likely landmark positions. Fig. 3 provides a conceptual illustration of shape subspace landmark refinement. A shape subspace $S$ is computed from a shape matrix $X$ consisting of landmark coordinates which we call a point set. Then, the similarity between the two subspaces is computed.
The key advantage of using shape subspace is its ability to represent the spatial relationships of a point set. while remaining invariant under affine transformations applied to the point set. This means that the shape subspace remains unchanged in case of positional shifts of the elbow bones in the images or size differences of the bone structures. Meanwhile, non-linear transformation of the point set, such as when the position of a point in the point set is shifted, alters the shape subspace. Although the order of the points in the shape matrices must be consistent [37], our approach always satisfies this condition because each point coordinate is obtained from each heatmap.
Figure 3: The process of shape subspace landmark refinement. Shape matrices $\mathbf{X}_{\text{train},i}$ and $\mathbf{X}_{\text{test},c_{j}}$ are derived from the landmark coordinates of the training data and the test heatmaps, respectively. Shape subspaces $\mathbf{S}_{\text{train},i}$ and $\mathbf{S}_{\text{test},c_{j}}$ are computed from the shape matrices, and the similarity between the two shape subspaces is compared. Note that, $z$ in $\mathbf{X}_{\text{train},i}$ and $\mathbf{X}_{\text{test},c_{j}}$ is set to zero to generate shape subspace from 2D coordinates.
2.3.1 Shape Subspace
We provide the mathematical definition of shape subspace. A shape subspace $\mathbf{S}$ is calculated from a $P\times 3$ shape matrix $\mathbf{X}$ that contains 3D coordinates of an object:
$\mathbf{X}=(\mathbf{r}_{1}\mathbf{r}_{2}\dots\mathbf{r}_{P})^{T}=\begin{pmatrix}x_{1}&y_{1}&z_{1}\\ \vdots&\vdots&\vdots\\ x_{P}&y_{P}&z_{P}\end{pmatrix},$ (3)
where $\mathbf{r}_{p}=(x_{p}\;y_{p}\;z_{p})^{T}$ for $1\leq p\leq P$ denotes the positional vector of the $p$–th point on an object.
Next, the shape matrix $\mathbf{X}$ is centered, and $\mathbf{X}_{c}$ is obtained to satisfy the relation $\sum_{p=1}^{P}\mathbf{r}_{p}=0$, making the shape subspace invariant to affine translation of the coordinates.
$\mathbf{X}_{c}=\mathbf{X}-\frac{1}{P}\sum_{p=1}^{P}\mathbf{r}_{p}.$ (4)
Then, Singular Value Decomposition (SVD) of the centered matrix $\mathbf{X}_{c}$ is computed, and we obtain left singular vectors $\mathbf{U}$, singular values $\mathbf{\Sigma}$, and right singular vectors $\mathbf{V}^{T}$. The basis vectors of shape subspace $\mathbf{S}$ are defined by the first $3$ columns of $\mathbf{U}$.

### v3

@@ Abstract
Ultrasound imaging of the medial elbow is crucial for early detection of Ulnar Collateral Ligament (UCL) injuries. Specifically, measuring the elbow joint space on ultrasound images enables the assessment of valgus instability. Automating this measurement requires a precisely annotated dataset to leverage recent advances in deep neural networks; however, no publicly available dataset has been proposed to date. In this work, we introduce a novel medial elbow ultrasound dataset with landmark annotations, facilitating the development of deep neural networks for joint space measurement. The dataset comprises 4,109 medial elbow ultrasound images from 22 participants, with precise landmark annotations on the humerus and ulna, performed under the supervision of three orthopedic surgeons. We evaluate several landmark detection models on our dataset and demonstrate its effectiveness in training models. Additionally, we propose Shape Subspace Refinement (SSR) to refine landmark heatmap peaks, improving the accuracy of heatmap-based landmark detection models. Furthermore, we show that the detected landmarks can be used for point-based segmentation of the humerus and ulna. The dataset is publicly available at https://github.com/Akahori000/Ultrasound-Medial-Elbow-Dataset.
Index Terms:
@@
Landmark detection, landmark refinement, shape subspace, medical diagnostic imaging, ultrasound dataset
[…]
=-21pt
@@
Figure 1: (a) The ultrasound images of the medial elbow obtained from different participants. (b) The enlarged views of the joint space positions. The left and right hyperechoic lines located in the middle of each image are the humerus and ulna, respectively. The space between humerus and ulna is the joint space to be measured.
I Introduction
Ultrasound imaging is an accessible, cost-effective, and noninvasive diagnostic technique widely used in clinical practice. In orthopedics, medial elbow ultrasound screening is commonly performed in athletes. Specifically, measuring the ulnohumeral joint space using ultrasound images provides critical insights into the risk of Ulnar Collateral Ligament (UCL) injury [1, 2]. An increase in the joint space under stress conditions suggests joint laxity and potential UCL damage of the elbow. Shanley et al. [2] measured joint space by defining landmarks on the edges of the ulna and humerus in ultrasound images of the medial elbow. Their study showed that pitchers with UCL injuries exhibit significantly greater joint space compared to those without injuries ($6.5\pm 1.2$ mm vs. $5.3\pm 1.2$ mm). To effectively measure joint laxity, Michinobu et al. [3] investigated the optimal elbow positioning during ultrasonography.
Several studies [4, 5, 6] have explored deep learning-based diagnosis using elbow ultrasound images. These studies have focused on osteochondritis dissecans or median nerve detection. However, none of the studies have discussed the automation of joint space measurements for diagnosing UCL injuries. Although precise and consistent measurement is required, it is challenging due to anatomical variability, noise, and artifacts. Fig. 1 illustrates ultrasound images of the medial elbow from different participants (a) and the enlarged views of the humerus and ulna near the joint space (b). These images show that bone shapes vary among individuals, and bone edges are often indistinct. Also, hyperechoic lines and artifacts often appear near bone surfaces or within the joint space.
@@ I Introduction
In this study, we introduce a medial elbow ultrasound dataset consisting of 4,109 images from 22 individuals. The dataset contains the precise landmark annotation on the humerus and ulna based on the expertise of three orthopedic surgeons. We evaluated joint space measurement methods on our proposed dataset using several landmark detection approaches. Although heatmap-based landmark detection methods generally achieve high accuracy, they may occasionally produce multiple peaks on a heatmap, leading to incorrect detection. To mitigate this, we propose Shape Subspace landmark Refinement (SSR) by measuring the geometric differences between detected and reference landmark positions. The results show that our measurement method is sufficiently accurate to diagnose UCL injuries. Finally, we demonstrate point-based segmentation of the humerus and ulna using detected landmarks as inputs, showing an application of our landmark detection in the medial elbow.
Figure 2: (a) The annotated landmarks: 1) Distal end of humeral trochlea, 2) Proximal end of ulnar coronoid process. Landmarks 1 and 2 are the joint space measurement points. 3) Tip of humeral medial epicondyle, 4) Deep end of the intersection of humeral medial epicondyle and ulnar collateral ligament, 5) Distal end of humeral medial epicondyle, 6) Proximal point adjacent to landmark 1 on the center of hyperechoic line of humeral trochlea, 7) Distal point adjacent to landmark 2 on the center of hyperechoic line of ulnar coronoid process, 8) Distal point of ulnar coronoid process. The white line on which landmarks 1 and 3 through 6 are located represents the humerus, while the white line on which landmarks 2, 7, and 8 are located represents the ulna. (b) The enlarged view of joint space. The yellow arrow is the joint space length to be measured. Landmarks 1 and 2 are the joint space measurement points. (c) The participant’s posture during the examination.
II Related work
II-A Landmark detection Methods
@@
Landmark detection, also known as keypoint detection, is a method of locating keypoints in images. Various methods have been proposed for pose estimation.
@@
Regression-based approaches [7, 8, 9, 10, 11, 12, 13] directly predict landmark coordinates. These methods are simple and computationally efficient, but often lack precision, especially in complex poses. Heatmap-based methods overcome this by generating heatmaps where the intensity represents the likelihood of a landmark’s location, achieving better spatial accuracy. Numerous studies have leveraged Convolutional Neural Networks (CNNs) for this purpose [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Among these, U-Net [24] and its extensions are widely used in localization tasks of medical images. The architecture utilizes skip connections to copy the feature maps from each stage of the encoder and directly pass them to the decoder’s corresponding stage. HRNet [25] maintains high-resolution representations throughout the whole process, allowing for accurate localization with multi-resolution features. Recently, Transformer-based heatmap models [26, 27, 28, 29] have been introduced, leveraging self-attention to model long-range dependencies. ViTPose [29] exemplifies this with a simple Vision Transformer architecture. Lastly, token-based methods [31, 30] learn the relationships between skeletal joints using tokens, where PCT [30] designs each token to represent a substructure of joints.
@@
Several landmark detection approaches have been applied to ultrasound images to measure anatomical structures [32, 33, 34, 35, 36]. U-Net and its extensions are used to measure left ventricle dimensions [33, 35]. Mask R-CNN-based architecture has been employed to evaluate developmental dysplasia of the hip from infant ultrasound images [34]. HRNet-based regression architecture is proposed to assess infant growth [36].
II-B Shape Subspace
@@
Shape Subspace is a method to represent the 3D points of a 3D structure as a linear subspace within a high-dimensional vector space. By measuring the similarity between two shape subspaces, the geometrical difference of 3D structures can be measured [37]. Shape subspace has been utilized to classify 3D objects and human faces [38, 39].
II-C Point-based Segmentation
@@
III Dataset Construction
III-A Data Collection
@@
We constructed an ultrasound medial elbow dataset of 4,109 images from 22 participants (ages 22–56, men and women, from five countries). All participants had no symptoms related to the elbow joint. The images were acquired using musculoskeletal ultrasonography (SONIMAGE MX1, KONICA MINOLTA, Tokyo, Japan) with an 11-MHz linear probe by two experienced orthopedic surgeons. The participants lay on a bed with their elbows flexed at 90 degrees as in [3]. Then, a surgeon moved the gel-applied probe around the medial elbow, as shown in Fig. 2(c). The images capture the medial epicondyle, UCL, medial surface of the humeral trochlea, and coronoid process of the ulna, a key region for diagnosing medial elbow conditions [3]. This procedure is the same as that used in routine medial elbow examinations (e.g., for baseball players). The image resolution is 528 × 672 pixels, corresponding to 0.0567 mm per pixel. In the images, the humerus appears on the left and the ulna on the right. Fig. 1 shows sample images of different participants with the enlarged joint space views. Representative images and additional information for each participant are provided in the supplementary material.
Figure 3: Visualization of shape subspace refinement (SSR). Shape matrices $\mathbf{X}_{\text{train},i}$ and $\mathbf{X}_{\text{test},c_{j}}$ are derived from the landmark coordinates of the training data and the test heatmaps, respectively. Shape subspaces $\mathbf{S}_{\text{train},i}$ and $\mathbf{S}_{\text{test},c_{j}}$ are computed from the shape matrices, and the similarity between the two shape subspaces is compared. Note that $z$ in $\mathbf{X}_{\text{train},i}$ and $\mathbf{X}_{\text{test},c_{j}}$ is set to zero to generate shape subspace from 2D coordinates.
III-B Annotation
@@
The obtained 4,109 images were manually annotated with landmarks at sub-pixel levels by the authors under the supervision of three orthopedic surgeons. The landmarks include two joint space measurement points: the distal end of the humeral trochlea, and the proximal end of the ulnar coronoid process, corresponding to the positions described in [2]. In addition to the joint space measurement points, six additional points on the humerus and ulna were annotated. These points are of particular interest to sonographers, who primarily focus on bone structures during ultrasound image interpretation. The annotated landmarks were meticulously reviewed and discussed by the three orthopedic experts, ensuring the reliability of the dataset.
@@
Fig. 2(a) shows the annotated landmarks: 1) Distal end of humeral trochlea, 2) Proximal end of ulnar coronoid process. Landmarks 1 and 2 are the joint space measurement points. 3) Tip of humeral medial epicondyle, 4) Deep end of the intersection of humeral medial epicondyle and ulnar collateral ligament, 5) Distal end of humeral medial epicondyle, 6) Proximal point adjacent to landmark 1 on the center of hyperechoic line of humeral trochlea, 7) Distal point adjacent to landmark 2 on the center of hyperechoic line of ulnar coronoid process, 8) Distal point of ulnar coronoid process. The yellow line segment between landmarks 1 and 2 in Fig. 2(b) shows the joint space to be measured.
III-C Ethical Approval and Informed Consent
This study was conducted in accordance with the principles of the Declaration of Helsinki. Ethical approval was obtained from the Ethics Committee of the Faculty of Medicine, University of Tsukuba (October 7, 2022, Approval No. 1517-3) and the Ethics Committee of the Center for Computational Science, University of Tsukuba (December 11, 2023, Approval No. 23-004; June 24, 2024, Approval No. 24-005). Informed consent was obtained from all participants involved in the study.
IV Method
IV-A Landmark Detection Models
@@
Several landmark detection models are compared: HRNet [25], ViTPose [29], and U-Net [24] as heatmap-based methods, YOLOv8 [13] as a regression-based method, and PCT [30] as a token-based method.
IV-A1 HRNet, ViTPose, and U-Net
@@
The landmark detection process of HRNet, ViTPose, and U-Net is represented in (1), where $f$ is the detection model and $\mathbf{x}\in\mathbb{R}^{W\times H\times 3}$ is the input image with width $W$, height $H$, and 3 channels. $f(\mathbf{x})$ outputs heatmaps of size $W^{\prime}\times H^{\prime}\times N_{K}$, where $W^{\prime}$ and $H^{\prime}$ are the heatmap width and height, and $N_{K}$ is the number of landmarks. Each heatmap corresponds to a specific landmark, meaning that $N_{K}$ heatmaps are generated. The $\arg\max(f(\mathbf{x}))$ operation returns the coarse landmark coordinates $\mathbf{p}^{\prime}\in\mathbb{R}^{N_{K}\times 2}$. The Unbiased Data Processing (UDP) function [41, 23] refines these to sub-pixel accuracy, yielding the final landmark positions $\mathbf{p}\in\mathbb{R}^{N_{K}\times 2}$, represented as $(x,y)$ coordinates of $N_{K}$ landmarks.
@@
IV-A2 PCT and YOLOv8
Equation (2) represents landmark detection using YOLOv8 and PCT. $g$ is the landmark detection model. The output $g(\mathbf{x})\in\mathbb{R}^{N_{K}\times 2}$ directly predicts the coordinates of the $N_{K}$ landmarks.
@@ IV-A2 PCT and YOLOv8
IV-B Proposed Shape Subspace Refinement
We propose Shape Subspace Refinement (SSR) to refine landmark positions in heatmap-based landmark detection. While heatmap-based models are trained to produce a single peak on each heatmap, they may occasionally generate multiple or unclear peaks. In such cases, measuring the geometrical similarities between detected and reference landmark positions using shape subspace can help select the most likely landmark positions. Fig. 3 provides a conceptual illustration of SSR. The shape subspaces for training and test samples $S_{train,i}$, $S_{test,c_{j}}$ are computed from their respective shape matrices $X_{train,i}$, $X_{test,c_{j}}$, which consist of landmark coordinates and are referred to as a point set. Then, the similarity between the two subspaces is computed.
The key advantage of using shape subspace is its ability to represent the spatial relationships of a point set while remaining invariant under affine transformations applied to the point set. This means that the shape subspace remains unchanged in case of positional shifts of the elbow bones in the images or size differences of the bone structures. Meanwhile, a non-linear transformation of the point set, such as when the position of a point in the point set is shifted, alters the shape subspace. Although the order of the points in the shape matrices must be consistent [37], our approach always satisfies this condition because each point coordinate is obtained from each heatmap.
IV-B1 Shape Subspace
The mathematical definition of shape subspace is provided below. A shape subspace $\mathbf{S}$ is calculated from a shape matrix $\mathbf{X}$ that contains 3D coordinates of an object:
$\mathbf{X}=(\mathbf{r}_{1}\mathbf{r}_{2}\dots\mathbf{r}_{K})^{T}=\begin{pmatrix}x_{1}&y_{1}&z_{1}\\ x_{2}&y_{2}&z_{2}\\ \vdots&\vdots&\vdots\\ x_{K}&y_{K}&z_{K}\end{pmatrix}$, (3)
where $\mathbf{r}_{i}=(x_{i}\;y_{i}\;z_{i})^{T}$ for $1\leq i\leq K$ denotes the positional vector of the ii–th point on an object.
Next, the shape matrix $\mathbf{X}$ is centered, and $\mathbf{X}_{c}$ is obtained to satisfy the relation $\sum_{i=1}^{K}\mathbf{r}_{i}=0$, making the shape subspace invariant to affine translation of the coordinates.
$\mathbf{X}_{c}=\mathbf{X}-\frac{1}{K}\sum_{i=1}^{K}\mathbf{r}_{i}.$ (4)
Then, Singular Value Decomposition (SVD) of the centered matrix $\mathbf{X}_{c}$ is computed, and left singular vectors $\mathbf{U}$, singular values $\mathbf{\Sigma}$, and right singular vectors $\mathbf{V}^{T}$ are obtained. The basis vectors of shape subspace $\mathbf{S}$ are defined by the first $3$ columns of $\mathbf{U}$ (Algorithm 1).